OpenStudy (anonymous):

Solve the equation on the interval 0 ≤ θ < 2π. 2 sin^2 θ = 3(cos θ + 1)

OpenStudy (anonymous):

So far I got Sin^2x = 3(cosx+1) /2

OpenStudy (anonymous):

Can Someone give me some basics about this because I feel lost

Parth (parthkohli):

\[\textbf{HINT}\]\[\cos^2\theta = {\cos\theta + 1 \over 2}\]

Parth (parthkohli):

So you have\[\cos ^2 \theta = 3\sin^2 \theta\]

OpenStudy (anonymous):

No, I will have\[\sin ^{2} = 3 \cos ^{2} x\]

Parth (parthkohli):

But the variable in your question is \(\theta\) :P